OpenStudy (anonymous):

72/3 reduced to lowest terms is what

OpenStudy (anonymous):

72 is divisible by 3

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

yep^^^^^^^^^

OpenStudy (anonymous):

new one 9/10*2 and2/3

OpenStudy (anonymous):

in simplest form

OpenStudy (anonymous):

these cannot be simplified

OpenStudy (anonymous):

\[{{9}\over{10}}\times{{2}{{2}\over{3}}} ={{9}\over{10}}\times{{{6+2}\over{3}}} \] \[={{9}\over{10}}\times{{{8}\over{3}}}\]

OpenStudy (anonymous):

now simplify it

OpenStudy (anonymous):

72/30

OpenStudy (anonymous):

again 72 and 30 are divisible by 3

OpenStudy (anonymous):

ty:}

OpenStudy (anonymous):

i ment 72/30 at te top(head slap)

OpenStudy (anonymous):

answer is \[{{72}\over{30}}= {{12}\over{5}}\]

OpenStudy (anonymous):

or 2.4

OpenStudy (phi):

Here is a trick you should learn. When you have a problem like \[ \frac{9}{10} \cdot \frac{8}{3} \] Notice that both 9 in the top and 3 in the bottom can be divided by 3. So do that immediately: \[ \frac{\cancel{9}3}{10} \cdot \frac{8}{\cancel{3}1} \] Also notice that 8 in the top and 10 in the bottom can be divided by 2 \[ \frac{3}{\cancel{10}5} \cdot \frac{\cancel{8}4}{1} \] Now multiply to get \[ \frac{12}{5} \] As a mixed number (divide by 5 and write the remainder as a fraction) \[ 2\frac{2}{5} \]